OpenStudy (anonymous):

Can someone explain to me the terminology between horizontal compression/expansion for this function: y=-f(.5x)-3

OpenStudy (anonymous):

I believe this is a horizontal expansion by a factor of 2, translated 3 units down and a reflection in the x-axis.

OpenStudy (experimentx):

check you last question.

OpenStudy (experimentx):

(.5x) <---- vertical compression if coefficient is greater than one, then it's expansion, if coefficient is less than one, then it's compression.

OpenStudy (anonymous):

I thought this function would be a horizontal expansion by a factor of 2..

OpenStudy (anonymous):

and vertical is outside the bracket.. o.O I am confused D:

OpenStudy (experimentx):

sorry ,,, my mistake .... didn't think carefully.

OpenStudy (anonymous):

am I correct? D:

OpenStudy (anonymous):

okay

OpenStudy (experimentx):

from here it seems, f(x) = 2 f(x) <--- is vertical expansion f(x) = f(2x) <---- is horizontal compression it seems, f(x) = 1/2 f(x) <--- is vertical compression f(x) = f(1/2 x) <---- is horizontal expansion

OpenStudy (anonymous):

yes, ohh, thank you! :)

OpenStudy (experimentx):

yw